OpenStudy (anonymous):

What interval includes all possible values of x, where –3(6 – 2x) ≥ 4x + 12?

OpenStudy (anonymous):

what are the answer choices?

OpenStudy (anonymous):

(–∞, –3] [–3, ∞) (–∞, 15] [15, ∞)

OpenStudy (anonymous):

solve for x : \[-3 (6-2x) \ge 4x + 12\] \[ -18 + 6x \ge 4x + 12\] \[2x \ge 30\] \[ x\ge 15\] So the answer is D ) 15 to +infinity

OpenStudy (anonymous):

do u have any idea?

OpenStudy (anonymous):

yep D

OpenStudy (anonymous):

@Ogziii is right

OpenStudy (anonymous):

it would be d